OpenStudy (anonymous):

A snow cone stand sells crushed, flavored ice in cone-shaped paper cups. The opening of the cup is 8 cm wide and the height of the cup is 12cm . What is the approximate volume of the cup itself?

OpenStudy (experimentx):

volume of cone = 1/3xarea of base x height

OpenStudy (anonymous):

is it 2000.96 thats what i got

OpenStudy (anonymous):

i mean 200.96

OpenStudy (anonymous):

depends on the flavor of the ice cream and the type of biscuit used for the cone volume of a cone is (1/3)Ah where A = the surface area of the base (here the opening) which is pi r^2. 8cm wide= radius 4 so 16pi times 12 = 603/3 = 201 cubic centimeters so yeah i get the same should be correct then ^^

OpenStudy (anonymous):

thx for the help

Directrix (directrix):

I ended up with 64 pi which gave 201.06. 1/3 pi 4^2 *12 = 64 pi